What is the weight/volume percent concentration of 7.5 grams glucose in 275 mL of solution?
Answer #1

mass of glucose = 7.5 g

volume of solution = 275 ml

w/v%= (mass of solute/ volume of solution)x 100 = (7.5/275)x 100=2.73 %